# Image-based Multimodal Models as Intruders: Transferable Multimodal Attacks on Video-based MLLMs

### v2

@@ 3.3 I2V-MLLM Attack
The proposed I2V-MLLM attack utilizes an IMM to produce adversarial video samples, targeting image-to-video cross-modal black-box attacks on V-MLLMs with significant transferability. By manipulating the intermediate features of vision models and projectors of IMMs, our approach generates adversarial video samples that interfere with the intermediate features of black-box V-MLLMs. The I2V-MLLM algorithm is illustrated in Appendix C, consists of three components: vision model attack, projector attack, and perturbation propagation.
3.3.1 Vision Model Attack
To enhance generalization in perturbing video features, I2V-MLLM disrupts both image features and spatiotemporal information extracted by the vision model. We first split the video $V$ into $K$ clips: $V=\{v^{1},v^{2},\ldots,v^{K}\}$, where $K=\text{total number of frames}\times\text{key-frame ratio}\,\beta$. We select the first frame $x^{k}$ from each clip $v^{k}$ as the key-frame, resulting in $K$ key-frames, $X=\{x^{1},x^{2},\ldots,x^{K}\}$, each capturing the essential information of their respective clips. Following [29], we extract spatiotemporal representations of $X$ using the vision model. This model independently encodes the $K$ frames, producing frame-level embeddings $F_{V}(X)\in\mathbb{R}^{K\times N\times D_{1}}$, where $F_{V}(\cdot)$ denotes the encoder of the vision model, $N$ is the number of patches per frame, and $D_{1}$ is the dimension of the embeddings. Frame-level embeddings are average-pooled along the temporal dimension to obtain a video-level temporal representation $F_{V}^{t}(X)\in\mathbb{R}^{N\times D_{1}}$, which implicitly incorporates temporal learning through the aggregation of $K$ frames. Similarly, the frame-level embeddings are average-pooled along the spatial dimension to obtain a video-level spatial representation $F_{V}^{s}(X)\in\mathbb{R}^{K\times D_{1}}$, which incorporate the spatial information of $K$ frames. The temporal and spatial features are concatenated to obtain the original video-level features $F_{V}^{ts}(X)=[F_{V}^{t}(X),F_{V}^{s}(X)]\in\mathbb{R}^{(N+K)\times D_{1}}.$ For the adversarial input $X_{adv}=\{x^{1}+\delta^{1},x^{2}+\delta^{2},\ldots,x^{K}+\delta^{K}\}$, we can similarly obtain the adversarial video-level features $F_{V}^{ts}(X_{adv})$. To disrupt the video-level features, I2V-MLLM optimizes the adversarial perturbations by minimizing the cosine similarity between the original and the adversarial video features:
@@ 3.3.1 Vision Model Attack
where $F_{V}^{ts}(X)_{i}$ and $F_{V}^{ts}(X_{adv})_{i}$ represent the $i$-th elements in the video-level features of the original and the adversarial video frames, respectively.
3.3.2 Projector Attack
To further disrupt V-MLLMs’ capacity for video-text multimodal tasks, I2V-MLLM interferes with the intermediate feature of the projector (e.g. Q-Former [20]), which plays an essential role in aligning visual and textual representations. We feed the projector with the original frame-level embeddings $F_{V}(X)$, the adversarial frame-level embeddings $F_{V}(X_{adv})$ from the vision model, and the caption set $T=\{t_{1},t_{2},\ldots,t_{M}\}$. After multimodal alignment, they are transformed into the original visual features $F_{P}^{v}(X)\in\mathbb{R}^{N_{1}\times D_{2}}$, the adversarial visual features $F_{P}^{v}(X_{adv})\in\mathbb{R}^{N_{1}\times D_{2}}$, and the textual features $F_{P}^{t}(T)\in\mathbb{R}^{N_{2}\times D_{2}}$. Here, $N_{1}$ and $N_{2}$ represent the number of visual features and the textual features, respectively. And $D_{2}$ denotes the dimension of these features. The captions are complete sentences generated based on the question $q$ and the answer $a$ using GPT-4o-mini [32]. For example, given the question $q$: ‘What is the man doing?’ and the answer $a$: ‘eat’, the corresponding caption $t$ would be: ‘The man is eating.’ To perturb the image features aligned with the text, I2V-MLLM optimizes the adversarial perturbations by minimizing the cosine similarity between the original and the adversarial visual features:
@@ 3.3.2 Projector Attack
3.3.3 Optimization and Perturbation Propagation
We optimize $\delta_{k}$ according to the following expression:
$\delta^{k}=\underset{\delta^{k}}{\arg\min}(\mathcal{L}_{total}),s.t.||\delta^{k}||_{\infty}\leq\epsilon,k=1,2,\ldots,K.$ (7)
Finally, we replicate $\delta^{k}$ to match the length of its corresponding video clip $v^{k}$, resulting in $\delta^{\prime k}$. We then construct the adversarial video by adding these perturbed clips to the original clips: $V_{adv}=V+\delta^{\prime}=\{v^{1}+\delta^{\prime 1},v^{2}+\delta^{\prime 2},\ldots,v^{K}+\delta^{\prime K}\}.$
4 Experiment
@@
Vanilla Chat-UniVi 52.48* 33.16* 1.98* 19.65 48.04 2.83 13.71 60.11 3.36 22.65 52.16 3.02 27.12
LLaVA-NeXT-Video 9.63 60.50 3.32 36.85* 38.05* 2.41* 13.71 59.97 3.37 26.51 48.93 2.88 21.67
VideoChat 8.22 61.50 3.38 15.86 51.11 2.98 35.03* 45.00* 2.69* 19.81 54.67 3.11 19.73
Video-LLaMA 11.93 58.88 3.26 31.82 41.01 2.52 14.40 59.33 3.34 63.96* 23.88* 1.72* 30.53
I2V CLIP-L/14 25.17 51.53 2.92 27.39 43.63 2.60 27.13 49.57 2.91 30.51 46.49 2.71 27.55
BLIP-2 43.39 38.72 2.33 40.54 33.84 2.21 63.09 26.08 1.82 74.91 17.07 1.39 55.48
I2V-MLLM InstructBLIP 36.74 43.29 2.56 37.61 36.17 2.30 54.26 31.99 2.10 69.90 20.58 1.58 49.63
MiniGPT-4 38.58 42.10 2.50 41.50 32.98 2.16 56.51 30.49 2.06 68.92 21.37 1.60 51.38
Table 2: The results on the MSVD-QA for VideoQA tasks. ASR (%) indicates attack success rate. Acc.(%) denotes the accuracy of the model’s predictions, while the Score represents GPT Score, which assesses the model and assigns a relative score to the predictions on a scale of 1 to 5. AASR represents the average ASR across all target models for each surrogate model. * indicates white-box attacks. A higher ASR or AASR indicates better adversarial transferability. The highest attack performance for each target model is shown in bold, and the second-highest in underline.
@@
Vanilla Chat-UniVi 55.10* 23.10* 1.68* 27.36 30.09 2.10 9.94 39.52 2.52 32.14 32.24 2.21 31.14
LLaVA-NeXT-Video 13.36 41.08 2.56 41.90* 24.25* 1.83* 8.50 39.56 2.53 35.23 30.41 2.13 24.75
VideoChat 11.88 41.35 2.57 25.40 30.72 2.14 27.47* 34.42* 2.29* 31.78 32.58 2.21 24.13
Video-LLaMA 13.71 40.78 2.55 37.88 26.14 1.91 8.73 39.53 2.52 63.66* 18.07* 1.53* 31.00
@@
BLIP-2 41.93 28.42 2.00 49.78 19.34 1.58 62.38 18.72 1.57 78.95 10.68 1.17 58.26
I2V-MLLM InstructBLIP 37.37 31.88 2.14 47.96 21.72 1.70 54.78 22.66 1.76 73.04 13.52 1.34 53.29
MiniGPT-4 38.60 30.94 2.11 49.47 21.12 1.67 56.41 21.95 1.73 73.28 13.63 1.32 54.44
Table 3: The results on the MSRVTT-QA for VideoQA tasks. The corresponding metrics and settings are consistent with those in Tab. 2.
@@ 4.1 Experimental setting
Datasets and models. Referring to the quantitative benchmarking framework proposed in [29], we evaluate our I2V-MLLM attack on VideoQA tasks using the validation set of MSRVTT-QA [41] and MSVD-QA [41], and on video understanding tasks using a subset of ActivityNet-200 [15]. We perform the proposed method on three IMMs: BLIP-2 [20], InstructBLIP [9] and MiniGPT-4 [46]. Our method is evaluated on four different V-MLLMs: Chat-UniVi [18], LLaVA-NeXT-Video [44], VideoChat [22], and Video-LLaMA [42], each with a Vicuna-7B [7] as the LLM.
Attack setting. In I2V-MLLM, we employ the projected gradient descent (PGD) [30] with a perturbation bound of $\epsilon=16$, an iteration number of $I=50$, and a step size of $\alpha=1$ for the attack process. The parameters $\lambda_{1}$ and $\lambda_{2}$ are both set to $1$, and the key-frame ratio $\beta$ is set to $30\%$. I2V attack, utilizing CLIP-L/14 [34] as the surrogate model, applies tailored perturbations to each frame of the video. For a fair comparison, the PGD parameters ($\epsilon=16$, $I=50$ and $\alpha=1$) in FMM, Vanilla, and I2V attacks maintain the same for our method. Additionally, in the FMM setup, the key-frame ratio $\beta$ is also set to $30\%$. All the experiments are conducted on a single NVIDIA-A6000 GPU.
Metrics. We use Attack Success Rate (ASR) to evaluate the effectiveness of adversarial examples on VideoQA tasks. It measures the percentage of successful attacks on questions the model answered correctly for clean videos. Answer correctness is evaluated using GPT-4o-mini [32], which checks whether the model’s prediction semantically aligns with the ground truth. We also provide the average ASR (AASR) across all evaluated V-MLLMs. A higher ASR or AASR indicates better adversarial transferability. To evaluate the model’s overall performance when encountering adversarial videos, we further employ GPT-assisted methods [29] to assess Accuracy (Acc.) and GPT-Score. Specifically, accuracy (Acc.) refers to the model’s prediction accuracy, while the GPT score (Score) assesses the quality of the model’s predictions, assigning a relative score on a scale from 1 to 5. GPT-4o-mini is used for evaluation due to its strong text understanding and cost efficiency. For detailed explanations of the metrics, see the Appendix B.
@@ 4.2 Attack performance
Evaluation of ASR. As shown in Tab. 2 and Tab. 3, I2V-MLLM achieves the best and near-best attack performance on the LLaVa-NeXT-Video, VideoChat, and Video-LLaMA. It attains ASR of 41.50%, 63.09%, and 74.91% for MSVD-QA, 49.78%, 62.38%, and 78.95% for MSRVTT-QA, respectively, even surpassing Vanilla and FMM attack methods in the white-box setting. For Chat-UniVi, I2V-MLLM achieves a suboptimal ASR of 43.39% for MSVD-QA and 41.93% for MSRVTT-QA, slightly below the Vanilla attack’s ASR of 52.48% for MSVD-QA and 55.10% for MSRVTT-QA in the white-box setting. These results demonstrate that I2V-MLLM performs better in the black-box setting, with its generated adversarial videos showing stronger cross-model transferability.
Alternatively, I2V-MLLM achieves the highest AASR compared to previous attack methods, achieving AASR of 55.48%, 49.63%, and 51.38% for MSVD-QA and 58.26%, 53.29%, and 54.44% for MSRVTT-QA when taking BLIP-2, InstructBLIP, MiniGPT-4 as surrogate models, respectively, significantly outperforming previous attack methods. This highlights that I2V-MLLM is capable of steadily generating high-quality adversarial video samples that deceive the unseen V-MLLMs.
Type Correct Detail Context Temporal Consistency
Clean 2.38 2.54 2.97 1.97 1.88
Vanilla 2.16 2.29 2.71 1.62 1.86
Vanilla* 2.07 2.25 2.64 1.55 1.74
I2V-MLLM 2.10 2.23 2.69 1.55 1.83
Table 4: The results on the ActivityNet-200 for video understanding tasks. All scores range from 1 to 5, with lower scores indicating better attack performance. * indicates a white-box attack. The highest attack performance is shown in bold, and the second-highest in underline.
V- MLLM $\lambda_{1}:\lambda_{2}$
1:1 1:2 1:3 2:1 3:1
Chat-UniVi 43.11 43.09 41.91 42.14 41.35
LLaVA-NeXT-Video 35.67 34.08 33.12 35.25 35.96
VideoChat 51.10 49.50 49.03 49.72 50.44
Video-LLaMA 64.57 64.11 62.98 63.56 61.94
AASR 48.61 47.69 46.76 47.67 47.42
Table 5: ASR (%) of the I2V-MLLM attack across different weight ratios of the vision model loss ($\lambda_{1}$) and projector loss ($\lambda_{2}$). A higher AASR indicates better adversarial transferability.
@@
As shown in Fig. 3, adversarial video samples generated from the I2V-MLLM attack cause V-MLLMs to produce responses that differ significantly from the clean answers, demonstrating that our method effectively misleads V-MLLMs and disrupts their ability to accurately interpret the video content.
4.3 Results on video understanding tasks
Video understanding tasks assess whether V-MLLMs have comprehended the content of a video by posing a range of questions about it. Following Maaz et al. [29], we use a subset of the ActivityNet-200 [15] dataset and employ GPT-4o-mini to evaluate the model’s responses to adversarial examples from five perspectives: Correctness, Detail Orientation, Contextual Understanding, Temporal Understanding, and Consistency. We compare our proposed I2V-MLLM attack with the Vanilla attack on LLaVA-NeXT-Video, using clean samples as a reference. Evaluations are performed on three attack types: white-box Vanilla attack, black-box Vanilla attack (using Video-LLaMA as a surrogate model), and I2V-MLLM attack (using BLIP-2 as a surrogate model).
As shown in Tab. 4, the black-box Vanilla attack induces minimal disruption, while the white-box Vanilla attack leads to the most significant interference. The I2V-MLLM performs similarly to the white-box Vanilla attack in terms of Correct, Detail, Context, and Temporal aspects, with only a drawback in Consistency. These results show that I2V-MLLM achieves comparable performance to the white-box Vanilla attack on video understanding tasks, further validating its effectiveness and transferability.
@@ 4.4 Ablation study
In this section, we provide ablation studies on the objective function, step size $\alpha$, number of iterations $I$, key-frame ratio $\beta$, and perturbation propagation in I2V-MLLM attack. Experiments are conducted on the MSVD-QA dataset for VideoQA tasks. We use BLIP-2 as the surrogate model and four different V-MLLMs as black-box models.
@@
Influence of loss functions and weight ratio. In Fig. 4, we provide ablation study on the components of the objective function used in our I2V-MLLM. The surrogate model is BLIP-2, and the generated adversarial videos are evaluated across four V-MLLMs. It can be observed that using either $\mathcal{L}_{V}$ or $\mathcal{L}_{P}$ alone achieves satisfactory attack performance. Combining both, which simultaneously disrupts low-level image features and the alignment between visual and textual modalities, further enhances the attack performance. We also vary the weights of the $\mathcal{L}_{V}$ and $\mathcal{L}_{P}$ to explore their relative relationship. As shown in Tab. 5, the AASR is highest when the ratio of $\lambda_{1}$ to $\lambda_{2}$ is 1:1. Therefore, we adopt this weight ratio in our experiments.
@@
Figure 4: ASR (%) of the I2V-MLLM attack with different loss functions.
@@
Influence of key-frame ratio and propagation. The number of key-frames used to optimize the perturbation, as well as the decision to propagate these perturbations across the entire video, significantly affects the transferability of adversarial video samples. Fig. 6 illustrates the results obtained with various key-frame ratios, comparing scenarios with and without perturbation propagation. It can be observed that as the key frame ratio increases, the generated adversarial samples show improved transferability. On the other hand, perturbation propagation substantially improves AASR by ensuring that all the frames extracted by unseen V-MLLMs are perturbed. As illustrated by the gain curve in the Fig. 6, the improvement from perturbation propagation initially rises with the key-frame ratio but then diminishes, reaching its maximum at 30%. With an AASR already high at a 30% key-frame ratio, further increases yield minimal gains, and perturbation propagation reaches its maximal benefit at this point. Therefore, we adopt a key-frame ratio of $\beta=30\%$. More experiments and analysis can be found in the Appendix B.

### v3

@@ 3.3 I2V-MLLM Attack
The proposed I2V-MLLM attack utilizes an I-MLLM to produce adversarial video samples, targeting image-to-video cross-modal black-box attacks on V-MLLMs with significant transferability. By manipulating the intermediate features of vision models and projectors of I-MLLMs, our approach generates adversarial video samples that interfere with the intermediate features of black-box V-MLLMs. The I2V-MLLM attack algorithm is illustrated in Appendix C, consists of three components: vision model attack, projector attack, and perturbation propagation.
Vision Model Attack
[…]
@@
where $F_{V}^{ts}(X)_{i}$ and $F_{V}^{ts}(X_{adv})_{i}$ represent the $i$-th elements in the spatiotemporal embeddings of the original and the adversarial video frames, respectively.
Projector Attack
To further disrupt V-MLLMs’ capacity for video-text multimodal tasks, I2V-MLLM interferes with the intermediate feature of the projector (e.g. Q-Former (li2023blip2bootstrappinglanguageimagepretraining)), which plays an essential role in aligning visual and textual representations. We feed the projector with the original frame-level embeddings $F_{V}(X)$, the adversarial frame-level embeddings $F_{V}(X_{adv})$ from the vision model, and the caption set $T=\{t_{1},t_{2},\ldots,t_{M}\}$. After multimodal alignment, they are transformed into the original visual features $F_{P}^{v}(X)\in\mathbb{R}^{N_{1}\times D_{2}}$, the adversarial visual features $F_{P}^{v}(X_{adv})\in\mathbb{R}^{N_{1}\times D_{2}}$, and the textual features $F_{P}^{t}(T)\in\mathbb{R}^{N_{2}\times D_{2}}$. Here, $N_{1}$ and $N_{2}$ represent the number of visual features and the textual features, respectively. And $D_{2}$ denotes the dimension of these features. The captions are complete sentences generated based on the question $q$ and the answer $a$ using GPT-4o-mini (openai2023gpt4omini). For example, given the question $q$: ‘What is the man doing?’ and the answer $a$: ‘eat’, the corresponding caption $t$ would be: ‘The man is eating.’ To perturb the image features aligned with the text, I2V-MLLM optimizes the adversarial perturbations by minimizing the cosine similarity between the original and the adversarial visual features:
@@
Vanilla Chat-UniVi 56.34* 30.40* 1.88* 12.64 48.60 2.87 6.00 59.05 3.35 21.78 53.53 3.06 24.19
LLaVA-NeXT-Video 9.33 59.05 3.30 52.45* 24.96* 1.53* 5.15 59.97 3.37 24.94 50.82 2.95 22.97
VideoChat 7.35 59.61 3.38 7.93 49.90 2.91 68.90* 23.81* 1.67* 20.74 54.15 3.09 26.23
Video-LLaMA 11.62 58.88 3.26 25.31 41.01 2.52 15.10 59.33 3.34 64.14* 23.88* 1.72* 29.04
I2V CLIP-L/14 32.17 51.53 2.92 33.39 43.63 2.60 41.13 49.57 2.91 36.51 46.49 2.71 35.80
BLIP-2 48.39 34.72 2.03 45.54 29.33 1.94 63.09 26.08 1.82 74.91 17.07 1.39 57.98
I2V-MLLM InstructBLIP 45.74 35.10 2.16 44.61 30.64 2.13 54.26 31.99 2.10 69.90 20.58 1.58 53.63
MiniGPT-4 43.58 37.02 2.21 46.50 27.98 1.76 56.51 30.49 2.06 68.92 21.37 1.60 53.88
Table 2: The results on the MSVD-QA for Zero-Shot VideoQA tasks. ASR (%) indicates attack success rate. Acc.(%) denotes the accuracy of the model’s predictions, while the Score represents GPT Score, which assesses the model and assigns a relative score to the predictions on a scale of 0 to 5. AASR represents the average ASR across all target models for each surrogate model. A higher AASR indicates better adversarial transferability. The best ASR for each target model under black-box attacks is highlighted in bold. * indicates white-box attacks for reference.
@@
Vanilla Chat-UniVi 52.19* 23.10* 1.68* 25.85 30.09 2.10 9.58 39.52 2.52 32.32 32.24 2.21 29.99
LLaVA-NeXT-Video 13.07 41.08 2.56 65.90* 15.06* 1.46* 8.50 39.56 2.53 35.23 30.41 2.13 30.68
VideoChat 12.64 41.35 2.57 25.27 30.72 2.14 63.46* 16.69* 1.42* 30.84 32.58 2.21 33.05
Video-LLaMA 14.33 40.78 2.55 37.91 26.14 1.91 8.83 39.53 2.52 64.11* 18.07* 1.53* 31.29
@@
BLIP-2 47.93 28.42 2.00 53.78 19.34 1.58 62.38 18.72 1.57 78.95 10.68 1.17 60.76
I2V-MLLM InstructBLIP 45.37 31.88 2.14 50.96 21.72 1.70 54.78 22.66 1.76 73.04 13.52 1.34 56.04
MiniGPT-4 46.60 30.94 2.11 49.47 21.12 1.67 56.41 21.95 1.73 73.28 13.63 1.32 56.44
Table 3: The results on the MSRVTT-QA. The corresponding metrics and settings are consistent with those in Tab. 2.
Optimization and Perturbation Propagation
We optimize $\delta_{k}$ according to the following expression:
[…]
Finally, we replicate $\delta^{k}$ to match the length of its corresponding video clip $v^{k}$, obtaining perturbed clip $\delta^{\prime k}$. The adversarial video is then constructed by applying pixel-wise addition to combine these perturbed clips with the original ones: $V_{adv}=V+\delta^{\prime}=\{v^{1}+\delta^{\prime 1},v^{2}+\delta^{\prime 2},\ldots,v^{K}+\delta^{\prime K}\}$. We term this approach Direct Propagation (DP). Due to the high similarity between consecutive frames, DP proves to be a simple yet effective method (see Sec. 4.3).
4 Experiment
@@ 4.1 Experimental setting
Datasets and models. Referring to the quantitative benchmarking framework proposed in (maaz2024videochatgptdetailedvideounderstanding), we evaluate our I2V-MLLM attack on Zero-Shot VideoQA tasks using the validation set of MSVD-QA and MSRVTT-QA. We perform the proposed method on three I-MLLMs: BLIP-2, InstructBLIP (dai2023instructblipgeneralpurposevisionlanguagemodels), and MiniGPT-4. Our method is evaluated on four different V-MLLMs: Chat-UniVi, LLaVA-NeXT-Video, VideoChat, and Video-LLaMA (zhang2023videollamainstructiontunedaudiovisuallanguage), each with a Vicuna-7B (vicuna2023) as the LLM.
Attack setting. In I2V-MLLM, we employ the projected gradient descent (PGD) (madry2019deeplearningmodelsresistant) with a perturbation bound of $\epsilon=16$, an iteration number of $I=50$, and a step size of $\alpha=1$ for the attack process. The parameters $\lambda_{1}$ and $\lambda_{2}$ are both set to $1$, and the key-frame ratio $\beta$ is set to $30\%$. I2V attack, utilizing CLIP-L/14 (radford2021learningtransferablevisualmodels) as the surrogate model, applies tailored perturbations to each frame of the video. For a fair comparison, the PGD parameters ($\epsilon=16$, $I=50$ and $\alpha=1$) in FMM, Vanilla, and I2V attacks maintain the same for our method. Additionally, in the FMM setup, the key-frame ratio $\beta$ is also set to $30\%$. All the experiments are conducted on an NVIDIA-A6000 GPU.
Metrics. We use Attack Success Rate (ASR) to evaluate the effectiveness of adversarial examples on Zero-Shot VideoQA tasks. It measures the percentage of successful attacks on questions the model answered correctly for clean videos. Answer correctness is evaluated using GPT-4o-mini, which checks whether the model’s prediction semantically aligns with the ground truth. We also provide the average ASR (AASR) across all evaluated V-MLLMs. A higher ASR or AASR indicates better adversarial transferability. To evaluate the model’s overall performance when encountering adversarial videos, we further employ GPT-assisted methods (maaz2024videochatgptdetailedvideounderstanding) to assess Accuracy (Acc.) and GPT-Score. Specifically, accuracy (Acc.) refers to the model’s prediction accuracy, while the GPT score (Score) assesses the quality of the model’s predictions, assigning a relative score on a scale from 0 to 5. GPT-4o-mini is used for evaluation due to its strong text understanding and cost efficiency. For detailed explanations of the metrics, see Appendix B.1.
@@ 4.2 Attack performance
Evaluation of ASR. As shown in Tab. 2 and Tab. 3, I2V-MLLM achieves the highest AASR compared to previous attack methods, achieving AASR of 57.98%, 53.63%, and 53.88% for MSVD-QA and 60.76%, 56.04%, and 56.44% for MSRVTT-QA when taking BLIP-2, InstructBLIP, MiniGPT-4 as surrogate models, respectively, significantly outperforming previous attack methods. I2V-MLLM (BLIP-2) achieves the best attack performance on Video-LLaMA and near-best attack performance on other target models. It achieves ASRs of 48.39%, 45.54%, 63.09%, and 74.91% on MSVD-QA, and 47.93%, 53.78%, 62.38%, and 78.95% on MSRVTT-QA, respectively, outperforming both the FMM and I2V attacks while achieving performance comparable to the white-box Vanilla attack.
@@ 4.3 Ablation study
In this section, we provide ablation studies on the loss functions, key-frame ratio $\beta$, perturbation propagation and different projectors in I2V-MLLM attack. Experiments are conducted on the MSVD-QA for Zero-Shot VideoQA tasks.
@@
Influence of loss functions. In Fig. 3, we provide ablation study on the components of the loss functions used in our I2V-MLLM. The surrogate model is BLIP-2, and the generated adversarial videos are evaluated across four V-MLLMs. It can be observed that using either $\mathcal{L}_{V}$ or $\mathcal{L}_{P}$ alone achieves satisfactory attack performance. Combining both, which simultaneously disrupts low-level image features and the alignment between visual and textual modalities, further enhances the attack performance.
@@
Figure 3: Attack success rates (ASR, %) of the I2V-MLLM attack with different loss functions.
@@
Influence of key-frame ratio and propagation method. The number of key-frames used to optimize the perturbation, as well as the decision to propagate these perturbations across the entire video, significantly affects the transferability of adversarial video samples. Fig. 4 illustrates the results obtained with various key-frame ratios, comparing scenarios with and without perturbation propagation. It can be observed that as the key frame ratio increases, the generated adversarial samples show improved transferability. Perturbation propagation substantially improves AASR by ensuring that all the frames extracted by unseen V-MLLMs are perturbed. As illustrated by the gain curve in the Fig. 4, the improvement from perturbation propagation initially rises with the key-frame ratio but then diminishes, reaching its maximum at 30%. With an AASR already high at a 30% key-frame ratio, further increases yield minimal gains, and perturbation propagation reaches its maximal benefit at this point. Therefore, we adopt a key-frame ratio of $\beta=30\%$.
Propagation Method / OFP BP DP
Chat-UniVi 37.69 34.03 44.63 48.39
LLaVA-NEXT-Video 23.34 32.07 39.72 45.54
VideoChat 23.67 42.26 54.60 63.09
Video-LLaMA 45.89 62.40 71.23 74.91
AASR 32.65 42.69 52.55 57.98
Table 4: Attack success rates (ASR, %) of the I2V-MLLM attack with different perturbation propagation method. ‘/’ represents no perturbation propagation. A higher AASR indicates better adversarial transferability.
Attack Surrogate Projector C-U L-N-V V-C V-L AASR
Vanilla C-U / 56.34* 12.64 6.00 21.78 24.19
L-N-V / 9.33 52.45* 5.15 24.94 22.97
V-C / 7.35 7.93 68.90* 20.74 26.23
V-L / 11.62 25.31 15.10 64.14* 29.04
LLaVA FC-Linear 41.31 43.32 50.31 56.72 47.92
mPLUG-Owl FC-Linear 44.97 42.07 49.77 57.16 46.74
I2V-MLLM BLIP-2 Q-Former 48.39 45.54 63.09 74.91 57.98
InstructBLIP Q-Former 45.74 44.61 54.26 69.90 53.63
MiniGPT-4 Q-Former 43.58 46.50 56.51 68.92 53.88
Table 5: Ablation Study on I-MLLMs with different projectors. Attack success rates (ASR, %) on the MSVD-QA validation set for Zero-Shot VideoQA tasks. * indicates a white-box attack. A higher AASR indicates better adversarial transferability. The highest attack performance for each target model in I2V-MLLM is shown in bold. Note: C-U: Chat-UniVi, L-N-V: LLaVA-NeXT-Video, V-C: VideoChat, V-L: Video-LLaMA.
Different perturbation propagation methods may yield varying results. We set the key-frame ratio $\beta=30\%$ and test three distinct perturbation propagation methods: Direct Propagation (DP), Optical Flow-based Propagation (dosovitskiy2015flownet) (OFP), and Bidirectional Linear Interpolation Propagation (dai2017deformable) (BP). The details of propagation methods are in Appendix B.1. As shown in Tab. 4, DP achieves the most significant improvement in AASR. Due to the high similarity between consecutive frames, DP proves to be a simple yet effective method. OFP suffers from added complexity and may distort the perturbation, resulting in lower effectiveness. BP’s slightly lower performance stems from its reliance on interpolation, which may dilute perturbation intensity compared to DP’s direct application. Therefore, we adopt DP in I2V-MLLM.
Influence of different projectors. The projectors of I-MLLMs are typically either FC-Linear (LLaVA (liu2023visualinstructiontuning), mPLUG-Owl (ye2023mplug)) or Q-Former (BLIP-2, InstructBLIP, MiniGPT-4). FC-Linear maps the visual features extracted by the vision encoder into the latent space of the LLM, whereas Q-Former further aligns visual and textual features before passing them to the LLM, enabling richer multimodal interactions. When Q-Former-based I-MLLM is employed as the surrogate, the projector attack module’s loss function is $\mathcal{L}_{P}=\mathcal{L}_{P_{v}}+\mathcal{L}_{P_{v2t}}$, where as for FC-Linear-based I-MLLM, it becomes $\mathcal{L}_{P}=\mathcal{L}_{P_{v}}$. As shown in Tab. 5, irrespective of the choice of surrogate model and projector, the adversarial transferability of I2V-MLLM significantly surpasses that of the vanilla attack. I-MLLMs equipped with Q-Former consistently outperform those with FC-Linear in generating transferable adversarial samples. This underscore the critical role of multimodal interactions in adversarial attacks on V-MLLMs. Therefore, in our main experiments, we exclusively employ Q-Former-based I-MLLMs as surrogate models to fully demonstrate the effectiveness of the I2V-MLLM approach.
Extended discussions on the influence of step size, iteration number, weights of loss functions and other experiments are detailed in Appendix B.2.